OpenStudy (tate_langdon):

Can someone check my work and show work Simplify the following expression. 14/15 - 3/10 A 3/5 B 11/5 C 19/30 D 6/5

OpenStudy (tate_langdon):

I say the answer is C

OpenStudy (tate_langdon):

@sunnnystrong

OpenStudy (tate_langdon):

@jhonyy9

OpenStudy (sunnnystrong):

\[\frac{ 14 }{ 15 }-\frac{ 3 }{ 10 }=\frac{ 28 }{ 30 }-\frac{ 9 }{ 30 }=\frac{ 19 }{ 30 }\] Yep good job:D @Tate_Langdon

jhonyy9 (jhonyy9):

yes you are right with C. sure